Lecture plan (preliminary version 18.08.23) TFY4345 Classical Mechanics autumn 2023

References: GPS = Goldstein, Poole, Safko 3. edition (2002) JAS = notater97.pdf LB = Linder, Brevik (Bookboon) KM = klassisk_mekanikk.pdf

Forelesninger:

Uke Tema  GPS JAS LB KM
34 Fundamental Principles 1 1 I 1

.

. Variational Principles and Lagrange’s Equations 2 2 II 2

.

. The Hamilton Equations of Motion  8  8 III 3
The Central Force Problem  3  3 IV 4
The Kinematics of Rigid Body Motion  4
The Rigid Body Equations of Motion  5
Oscillations  6
The Special Theory of Relativity  7
Canonical Transformations  9
Hamilton–Jacobi Theory  10

First lecture: Monday 21. November. (Week 34.) Final lecture: Wednesday 22. November. (Week 47.) Q & A before exam: To be agreed. Exam: 27.11.23, 9-13.